The exo planet closest to us currently (as of 2019) Proxima b orbits its star Proxima Centuri at an average distance of 7.5 ×10^9 m. If the average gravitational potential energy of this star-planet system is -1.7 × 10^34 J, what is the magnitude of the approximate gravitational force exerted on this planet by the star? Hint: you do not need the masses of this star or the plant to answer this question
Answer #1

given

r = 7.5*10^9 m
U = -1.7*10^34 J

let m1 is the mass of the star and m2 is the mass of the planet.

we know, the average gravitational potential energy of this star-planet system, U = -G*m1*m2/r

force between the star and the planet,
F = G*m1*m2/r^2

= -(-G*m1*m2/r)*(1/r)

= -U/r

= -(-1.7*10^34)/(7.5*10^9)

= 2.27*10^24 N <<<<<<<<---------------Answer